Describe and analyze the following data, calculating elasticities. For a product, when price increases from $10 to $12, the demand increases from 200 to 225 [Calculate elasticity]; On the other hand when price increases from $ 16 to $ 18, the demand decreases from 230 to 225 [Calculate elasticity]. Explain what seems to be happening.
Answer #1

Answer

Elasticity of demand=(change in quantity/average quantity)/(change in price/average price)
Change in quantity=225-200=25
average quantity=(225+200)/2=212.5
change in price=12-10=2
average price=(12+10)/2=11
The elasticity of demand=(25/212.5)/(2/11)

=0.647058824=0.64

the elasticity is 0.64 it is positive means the demand curve is upward sloping so the good is a Giffen good.

Elasticity of demand=(change in quantity/average quantity)/(change in price/average price)
Change in quantity=225-230=-5
average quantity=(225+230)/2=227.5
change in price=18-16=2
average price=(18+16)/2=17
Elasticity of demand=(-5/227.5)/(2/17)

=-0.186813187

the good is inelastic as the demand elastcity is below 1 in absolute value. and the good is neccesity.

At the lesser price, the good is Giffen good but at higher prices the good is a necessity because as price increases the relative income decreases and the demand for the good is more necessary than just a poor quality option.

.